what is the slope and y-intercept for 3y + 3 - 2y - 5x/2 =0
Answer by mananth(16946)

3y + 3 - 2y - 5x/2 =0
y=5x/2-3
compare with the equation y= mx+b
m= slope = 5/2
b= y intercept = -3